OpenStudy (anonymous):

How to calculate the inverse function of f(x) = (exp(x))+2)/(exp(-x))?

OpenStudy (fibonaccichick666):

so first, let me rewrite: \[f(x)=\frac{e^x+2}{e^{-x}}\] Correct?

OpenStudy (anonymous):

yes that is correct :)

OpenStudy (fibonaccichick666):

ok, so now, in order to find the inverse, let's first talk about what the inverse is. The inverse "undo's" a function, but it also happens to be what happens when you swap the domain and range. So do you have any ideas for our first step?

OpenStudy (anonymous):

Can you tell why the inverse exists?

OpenStudy (anonymous):

Like you said, we exchange the domain and the range, so that's what I tried to do. So, we "swap" x for y and y for x and rearrange

OpenStudy (anonymous):

where f(x) = y

OpenStudy (anonymous):

then, first I wrote \[y = e ^{2x} + 2e ^{x}\]

OpenStudy (fibonaccichick666):

Yup but before you do that, graph the function, as Abhishek pointed out, there does not have to be an inverse

OpenStudy (fibonaccichick666):

hold off on that

OpenStudy (anonymous):

I don't quite understand what Abhishek69 meant by that

OpenStudy (fibonaccichick666):

technically* you should test for an inverse first, I just skipped the step because I knew the answer. My bad so in order for a fn to have an inverse, it must be 1-1 or pass the horizontal line test

OpenStudy (fibonaccichick666):

aka for every x, there is a unique y

OpenStudy (anonymous):

yep okay

OpenStudy (fibonaccichick666):

does this eq pass?

OpenStudy (anonymous):

yes it does

OpenStudy (fibonaccichick666):

cool ok so now, your rewrite was correct, continue down the path

OpenStudy (anonymous):

I then tried to take the natural log of both sides, and do the usual stuff (hopefully correct anyway). But in the end I got something that was quite different to the answer (it was a demonstration in the lecture I tried to redo). The answer I got was \[\frac{ 1 }{ 3 } \ln(\frac{ y }{ 2 })\]

OpenStudy (fibonaccichick666):

so you made a fundamental error, ln(x+y) cannot be simplified at all

OpenStudy (fibonaccichick666):

let's walk through what you did

OpenStudy (anonymous):

so, \[\ln(y) = \ln(e ^{2x}) \times \ln(2e ^{x})\]

OpenStudy (anonymous):

is that correct? I took the log of the whole of the right-hand side

OpenStudy (fibonaccichick666):

no

OpenStudy (anonymous):

ah.

OpenStudy (fibonaccichick666):

they are not multiiplied

OpenStudy (anonymous):

yes but i used the log rule

OpenStudy (fibonaccichick666):

they are addition,

OpenStudy (fibonaccichick666):

again ln(x+y) cannot be simplified

OpenStudy (fibonaccichick666):

ln(xy)=(lnx)+(lny)

OpenStudy (fibonaccichick666):

there is no rule for this situation

OpenStudy (fibonaccichick666):

so rewrite from the beginning distribution you did

OpenStudy (anonymous):

okay

OpenStudy (fibonaccichick666):

SO i'll give you a hint. Your first step is going to be: complete the _____.

OpenStudy (anonymous):

wait wait wait wait

OpenStudy (fibonaccichick666):

hm?

OpenStudy (anonymous):

I was being an idiot. I didn't do what I did, that was the "fail" attempt that i forgot to cross out

OpenStudy (fibonaccichick666):

o.O ok?

OpenStudy (anonymous):

yeah anyway, if i now take the log of each term, that's what i'd get to in the end. (with the 1/3 blah answer). now I complete the square?

OpenStudy (fibonaccichick666):

no

OpenStudy (fibonaccichick666):

there is no 1/3

OpenStudy (anonymous):

you're right, it's wrong nonetheless

OpenStudy (anonymous):

i was going to say, now if i let x = exp(x)

OpenStudy (fibonaccichick666):

so let's start here \[y=e^2x+2e^x\]

OpenStudy (anonymous):

and write y = x^2 + 2x, would that do something?

OpenStudy (fibonaccichick666):

no

OpenStudy (fibonaccichick666):

so let's start here \[y=e^{2x}+2e^ x\]

OpenStudy (fibonaccichick666):

just trust me ok, i'll walk you through

OpenStudy (fibonaccichick666):

first step?

OpenStudy (anonymous):

really not a clue, pardon

OpenStudy (fibonaccichick666):

ok, so remember, swithc the variables to start

OpenStudy (anonymous):

okay, so now we swap y and x?

OpenStudy (fibonaccichick666):

yep type that in, you can copy and paste my LaTex commands using right click show math as

OpenStudy (anonymous):

oh. I think this is it. I forgot to do the swapping from the start.

OpenStudy (fibonaccichick666):

nope, please... please... I'm tired and losing patience, just do step by step with me please?

OpenStudy (anonymous):

But I'm sorry, I have a lecture to go to in half an hour so I need to get ready. I'll try this question again later. Thank you for your time, and I'm sorry about this!

OpenStudy (fibonaccichick666):

wait I have the answer pleasssseeee

OpenStudy (fibonaccichick666):

it's just messy if you do it with me we can be done in 10 min or less

OpenStudy (anonymous):

I would love to but I really gotta dash :/ Thank you very much for the offer though

OpenStudy (fibonaccichick666):

ugh, ok

OpenStudy (fibonaccichick666):

just complete the square then simplify then ln then simplify then exp then simplify then ln again and done

OpenStudy (anonymous):

okay okay, thanks a lot! will try again later after physics. have a nice day :)

OpenStudy (fibonaccichick666):

you too. took me 8 steps, so when you do just try ok?

OpenStudy (anonymous):

yep